OpenStudy (anonymous):

Find the exact value of the trigonometric function at the given real number. 1. cos (-13pi/20) 2. cos (10pi/3)

OpenStudy (e.mccormick):

Just to confirm: (-13pi/20) or (-13pi/2) ? I ask because the second is far simpler than the first.

OpenStudy (e.mccormick):

And with 2. cos (10pi/3)....

OpenStudy (anonymous):

i just need an explanation on how to solve the second one

OpenStudy (e.mccormick):

2. cos (10pi/3) ? Do you know what coterminal angles are?

OpenStudy (anonymous):

isn't it the angle that's given is the terminal and the angle around is the other one?

OpenStudy (e.mccormick):

Well, on the unit circle there are many angles that represent the same thing. The first of these you see is 0. \(0=2\pi\). You can extend this and say: \(0=2\pi = 4\pi = 6\pi\) and so on.

OpenStudy (anonymous):

so is 10pi/3 the same as another angle on the circle?

OpenStudy (e.mccormick):

Exactly! The reason why this works is because the results of sine, cosine, tangent, and so on of those values are the same. So, what is a coterminal of \(\frac{10\pi}{3}\) that is between 0 and \(2\pi\)

OpenStudy (e.mccormick):

Subtract \(2\pi\) from \(\frac{10\pi}{3}\). But be careful to make it into a fraction over 3 to do this!

OpenStudy (anonymous):

4pi/3? and why do you have to subtract 2pi from 10pi/3?

OpenStudy (e.mccormick):

It would be a full circle less, which is \(2\pi\). And yes, that makes \(\frac{4\pi}{3}\) a cotermnal you could use. Let me draw why.

OpenStudy (e.mccormick):

You see how it adds a recolution?